Calculate the concentration of all species in a 0.230 M M C6H5NH3Cl C 6 H 5 N H 3 C l solution. Enter your answers numerically separated by commas. Express your answer using two significant figures.

[C6H5NH3+], [Cl−], [C6H5NH2], [H3O+], [OH−]
Answer #1

C6H5NH3Cl (aq) + H2O (l) C6H5NH2 (aq) + Cl- (aq) + H3O+ (aq)

So, [H3O+] = [Cl-] = [C6H5NH2]

C6H5NH3Cl is salt of weak base (C6H5NH2) and strong acid (HCl).

Pkb of aniline = 9.4

pH of salt solution of weak base strong acid is

pH = 7 - ( pKb - logCsalt)

Given concentration of salt (Csalt) = 0.230 M.

Then,

pH = 7 - {9.4 - log(0.230)}

= 7 - ( 9.4 + 0.64)

= 7 - 5.0

= 2.0

Then, [H3O+] = 10-2.0 = 0.01 M

[OH-] = Kw/[H3O+]

[OH-] = (10-14/ 0.01) = 1.0 × 10-12 M

[H3O+] = [Cl-] = [C6H5NH2] = 0.01 M.

[C6H5NH3+] = 0.23 M